o:

Evaluate the following expression using the values given: Find −a2 − 3b3 + c2 + 2b3 − c2 if a = 3, b = 2, and c = −3. (1 point) −67 −17 −31 1

Shadow:

So you want me to be a calculator? ._.

Shadow:

You inputted the variables in, you know exactly what you're doing Dora

o:

The answers given do not match up with the calculator

Shadow:

Gimme a min, almost home

o:

Okay

o:

i edited the question so it's in it's original state

BlankSpace:

The inputted variables have been reset

BlankSpace:

brb changing the exponents

BlankSpace:

Evaluate the following expression using the values given: Find −a^2 − 3b^3 + c^2 + 2b^3 − c^2 if a = 3, b = 2, and c = −3. (1 point) −67 −17 −31 1

Shadow:

-.-

Shadow:

Dora, you understand that exponents are important, right?

BlankSpace:

I know, I accidentally forgot the thingy

Shadow:

\[−a^2 − 3b^3 + c^2 + 2b^3 − c^2 \] \[a = 3, b = 2, c = −3\] \[-(3)^2 - 3(2)^3 + (-3)^2 + 2(2)^3 - (-3)^2\]

BlankSpace:

Ok thank you Shads

BlankSpace:

I did the rest ^.^

Shadow:

You did the math?

BlankSpace:

well duh. im not gonna use a calculator

Shadow:

What did you get?

BlankSpace:

only on square roots

BlankSpace:

-17

Shadow:

Correct, good job

BlankSpace:

Thanks Shads